OpenStudy (anonymous):

Solve for x: the quantity of x plus 4 over 3 = 2 ?

OpenStudy (anonymous):

(x+4)/3=2 is this ur question?

OpenStudy (anonymous):

@susan3 yes

OpenStudy (anonymous):

x+4=6 x=6-4 x=2

OpenStudy (anonymous):

@susan3 thank you SO much

OpenStudy (anonymous):

no probelm @InstagramDoll